We’ve established that heights of 10-year-old boys vary according to a Normal distribution with μ = 140 cm and σ = 5 cm.

What proportion is between 150 and 140 cm?

a.

68%

b.

50%

c.

97.72%

d.

47.72%
Answer #1

Solution :

Given that ,

mean = \mu = 140

standard deviation = \sigma = 5

P(140< x <150 ) = P[(140 - 140) / 5< (x - \mu ) / \sigma< (150 - 140) / 5)]

= P( 0< Z < 2)

= P(Z < 2) - P(Z < 0)

Using z table   

= 0.9772 - 0.5

=0.4772

=47.72%